# The Rise of Small Language Models in Healthcare: A Comprehensive Survey

### v1

@@ 5.1. Current state-of-the-art
Discussion on comparative analysis of SLMs: SLMs have rapidly evolved for a diverse range of healthcare applications. Table 6 summarizes existing healthcare SLMs, detailing their base architectures, adaptation techniques, and target NLP tasks. Most SLMs are built on lightweight architectures like GPT-2, Phi-2, Mamba, Qwen, and LLaMA-2, often adapted with techniques such as fine-tuning and in-context learning. However, fewer models leverage quantization. Models like BioMistral, BioMistral-NLU, MentalQLM, and PMC-LLAMA exemplify state-of-the-art performance through multi-task adaptability, multilingual benchmarking, and specialized reasoning tasks (e.g., CoT QA). Yet, many efforts are still concentrated on QA and CLS, while areas like clinical SUM, temporal reasoning, and wellness dimension extraction remain underexplored. Additionally, we mapped the availability of SLMs within healthcare domain.
At present, healthcare models employ knowledge-aware methods, including ontology grounding using UMLS or SNOMED CT and RAG to enhance factual correctness. Integrating advanced interpretable reasoning mechanisms, such as metacognition, holds promise for enhancing the trustworthiness and reliability of clinical decision support systems. There is a pressing need to broaden the scope of tasks beyond traditional disease-centric applications like question answering and classification, toward holistic patient care, including the extraction and interpretation of emotional, social, behavioral, and wellness dimensions. Establishing standardized benchmarks for emerging clinical NLP tasks - wellness assessment, mental health triage, and personalized care planning will be vital to evaluating real-world utility and scientific progress. Moreover, Magnini et al. (2025) evaluates the use of locally deployed open-source generic SLMs for medical chatbots, focusing on hypertension self-management to ensure privacy and reliability. It demonstrates that lightweight SLMs can perform competitively in key tasks, offering a viable, privacy-preserving alternative to large proprietary models.
@@ 5.2. The role of SLM in healthcare
Beyond the performance of standalone models, we explore the role of SLMs in synergy with LLMs, emphasizing their complementary strengths, capabilities, and applications in healthcare research. The potential of collaborative and competitive roles of SLMs and LLMs is underappreciated in fostering efficient and accessible AI models for clinical decision making. Amid the broad spectrum of SLM applications observed in prior studies on general models (Chen and Varoquaux, 2024), this section delineates their specific relevance to NLP-driven decision-making tasks within the scope of healthcare informatics.
@@ 5.3. Assessment methods
Accuracy and Judgment: Chen et al. (2023) develop a novel, physician-curated adversarial question dataset rooted in real-world clinical scenarios, along with a comprehensive 17-metric evaluation framework designed to assess alignment and contextual relevance to clinical practice. Selene Mini is an open-source SLM designed for high-accuracy evaluation across diverse tasks [^14] https://huggingface.co/blog/AtlaAI/selene-1-mini. Trained with a principled blend of supervised and preference optimization on curated data, it outperforms 11 benchmarks of SLMs as a judge, demonstrating robust, domain-adaptable, and format-resistant performance for real-world deployment.
@@
Trustworthiness: Benchmarking LLMs in healthcare requires standardized, domain-specific metrics to ensure safety, accuracy, and ethical integrity while addressing data privacy, bias, and explainability (Budler et al., 2025; Aljohani et al., 2025). Truthfulness is critical in language models to prevent misinformation. Using the TruthfulQA benchmark, BioMistral-7B shows strong performance, outperforming GPT-3.5 Turbo by 4%, though no model excels across all categories, highlighting task-specific strengths and limitations (Labrak et al., 2024). Faithfulness in AI-generated medical information remains a critical concern, and recent efforts focus on understanding its causes, developing evaluation metrics, and proposing mitigation strategies to ensure factual accuracy across generative medical AI applications (Xie et al., 2023b).
Memory and energy efficiency: Energy efficiency is a high priority for SLMs, especially when used on battery-powered devices. MELODI is a novel framework and dataset designed to monitor and analyze energy consumption during LLM inference, revealing significant efficiency disparities and offering a foundation for sustainable, energy-conscious AI deployment (Husom et al., 2024). Response length significantly impact power consumption, and concise outputs can reduce energy usage. We measure the reduced environmental impact of healthcare SLM development through carbon footprints in Table 5.3. Greenhouse equivalences can be calculated with a well-establish approach [^16] https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator.
@@ 6. Discussion and conclusion
Our goal is to provide healthcare professionals with practical insights on deploying SLMs in real-time applications, rather than explaining their internal workings. SLMs assist healthcare professionals by extracting key clinical concepts—such as social determinants of health, quality of life, and therapeutic interventions—from unstructured text. They support clinical decision-making through automated diagnosis assistance, personalized treatment recommendations, and summarization of lengthy patient records. Additionally, SLMs enhance real-time clinical decision support by identifying rare diseases, drug interactions, and aligning patient data with medical literature. SLMs support clinical research by structuring EHR data and monitoring adverse events, thereby reducing manual chart reviews and enhancing data usability. They automate clinical coding and billing by mapping free-text entries to standardized codes, minimizing administrative burden and claim errors. SLMs also enable clinical QA by extracting relevant insights from research literature and patient records, accelerating evidence-based investigations. SLMs have the potential to enhance patient-centered care by analyzing sentiment and monitoring mental health to detect risks like worsening depression. They empower virtual health assistants and triage systems, facilitating personalized engagement and real-time conversational support. Additionally, SLMs improve healthcare accessibility through Plain Language Adaptation (PLA) - simplifying complex medical information for individuals with limited health literacy.
@@ 6.1. Challenges and research directions
The contrast between the theoretical capabilities and real world limitations is a fascinating space to explore. SLMs in NLP-driven healthcare informatics are facing a significant benchmarking challenge. Unlike general-domain NLP, healthcare informatics relies heavily on proprietary, heterogeneous data, limiting reproducibility and transparency. Public datasets offer partial solutions but often lack representativeness. As models grow in complexity and autonomy, evaluation frameworks must evolve to incorporate safety, ethics, and trustworthiness. Standardized benchmarks are essential to ensure consistent assessment, with key criteria including “doctor-like" communication—demonstrating medical expertise, sound reasoning, and clinically relevant diagnostic insight. Progress requires collaborative efforts, sustainable funding, regulatory alignment, and a workforce skilled in both informatics and health sciences. Edge deployment is critical for expanding access to AI tools in underserved regions, enabling point-of-care decision support, and reducing reliance on cloud-based APIs that may raise data privacy concerns. Interoperability is the model’s ability to seamlessly integrate and operate within the existing Health Information Exchanges (HIEs) from different sites. Another significant advantage of SLMs is their potential when used in multi-agent AI systems—collections of specialized agents that collaborate to solve complex tasks. Instead of relying on a single monolithic model like an LLM, multi-agent systems distribute tasks among several domain-specific SLMs. Future research must focus on developing modular interfaces, shared embedding spaces, and universal chaining frameworks that enable plug-and-play interoperability across models of varying sizes and capabilities. We posit the SLM development as a multi-dimensional solution for advancing NLP-centered healthcare informatics. Several active research directions in the development of generic SLMs are being effectively adapted to advance healthcare-specific SLM development. We enlist significant developments in healthcare as follows.
@@
Human-in-the-loop systems integrate SLMs, LLMs, and expert oversight to ensure safe and effective deployment. SLMs assist with triage and uncertainty flagging, while LLMs manage complex synthesis. Embedding human feedback enhances trust and supports continual model refinement. Advancing flexible architectures, evaluation frameworks, and regulatory standards is essential, with future research focusing on efficient evaluators for clinical tasks, emphasizing factuality, safety, and uncertainty (Han et al., 2024; Xie et al., 2023b; Budler et al., 2025). Reducing manual chart review. As LLMs are preferentially aligned with human-preferred responses and surpass human capabilities, the feasibility and reliability of human evaluation diminish, introducing a weak-to-strong generalization, wherein smaller, less capable models supervise and guide the alignment of more powerful models (Burns et al., 2023). This paradigm has potential to reduce manual chart review efforts, enabling scalable, cost-effective, safe and factual preference alignment that exceed clinician expertise. Summarizing medical reports. SLMs perform high-quality abstractive summarization by distilling structured rationales from LLMs (Wang et al., 2025a). These rationales encapsulate important structural and semantic cues from the document, broken into high-level concepts, text-spans that support those themes, and relationships among key entities in the document. These structured rationales serve as interpretable reasoning traces that guide summary generation, similar to CoT prompting but in a structured, modular format. SLMs are rationale-aware, lightweight summarizers that are trained using intermediate reasoning signals setting a new paradigm for transparent and cost-effective summarization. Bias detection and de-biasing models. Because SLMs are often fine-tuned on relatively small clinical corpora (e.g., institutional EHRs or specialty-specific datasets), they are at high risk of inheriting data biases or overfitting to local linguistic patterns (Sapkota et al., 2025). This can manifest in poor generalization to diverse patient populations, inaccurate risk stratification, or propagation of health disparities in AI-driven decision-making. Future research must incorporate domain-aware debiasing techniques such as adversarial training, fairness-aware objective functions, and counterfactual data augmentation. Rigorous bias audits and alignment assessments should become standard in the development lifecycle of healthcare SLMs. Automated Machine Learning (AutoML). While existing model compression techniques have shown significant progress, they continue to rely heavily on manual design—particularly in defining suitable student architectures for knowledge distillation, which demands considerable human effort. The application of AutoML to LLM remains limited due to the significant computational burden involved in scaling down models with billions of parameters. SLMs address this challenge by evaluating architectural components like model depth and attention head width within a reduced parameter space (Liu et al., 2024e). Shen et al. (2024) reduce the search complexity by incorporating optimized initialization strategies, accelerating convergence. Therefore, the models must integrate AutoML approaches—meta-learning and neural architecture search, that offer a promising direction to mitigate this dependency. This will enhance compression efficiency without compromising performance.
@@ 6.2. Impact of SLMs
SLMs offer infrastructure-friendly deployment on edge devices using optimized hardware such as NVIDIA Jetson or Qualcomm AI accelerators. Software optimization techniques (distillation, quantization and pruning) significantly lower computational demands, enabling model adoption without costly hardware investments. However, challenges such as frequent model updates and deployment across heterogeneous devices may introduce new scalability and maintenance complexities. Given these considerations, investigating the impact of model development is necessary to support responsible application. Environmental and economic trade-offs: Training GPT-3 with 175 billion parameters consumed 1287 MWh of electricity, and resulted in carbon emissions of 502 metric tons of carbon, equivalent to driving 112 gasoline powered cars for a year (Patterson et al., 2021). The environmental impact of AI is an escalating concern (Rillig et al., 2023) where SLMs offer a partial solution by reducing the training and inference energy consumption by up to 80% (Liu and Yin, 2024). This reduction directly translates into a smaller carbon footprint, making SLMs a more sustainable choice as evident from Table 5.3. This efficiency is valuable in healthcare, aligning with the growing movement toward green AI, where the need of real-time solutions must be balanced with sustainability goals. The future directions for environmentally conscious green (small/large) language model development in healthcare promotes compression techniques discussed in §4. Ethical and Social Impact. The integration of SLMs into clinical workflows alleviates provider burden while promoting equitable access, reducing digital disparities, and enhancing patient-centered care. Their lightweight nature enables deployment in low-resource settings, including rural clinics, mobile health units, and edge devices, thereby extending AI-driven support to underserved populations. By lowering the technical and financial barriers associated with large models, SLMs democratize the use of advanced NLP tools across diverse healthcare institutions (Mittal et al., 2024). Additionally, SLMs support data de-identification by enabling efficient, on-device removal of personally identifiable information (PII) from clinical texts, facilitating privacy-preserving data sharing and compliance with healthcare regulations. SLMs can support patient friendly adoption through plain language adaptation, and empower patients with timely, comprehensible information. However, addressing fairness, bias, and equitable access remains critical (Sapkota et al., 2025). While the narrower focus of model training datasets allows for more effective bias filtering, it also risks excluding underrepresented groups or contexts. Therefore, careful dataset design is essential to ensure inclusive and responsible AI deployment. Synthesis and Perspectives. The goal of SLMs is to prioritize efficiency and accessibility. While, they mitigate environmental, economic, and ethical challenges associated with LLMs, the success of these methods largely depend on interdisciplinary collaborations, for example, hardware engineers optimizing for energy efficiency, clinicians validating medical tools, and policymakers ensuring equitable access. Future work in this direction requires standardized benchmarks for fairness and environmental impact, a step toward sustainable, human-centric AI.
@@ 6.3. Concluding remarks
This work underscores the transformative potential of SLMs in democratizing AI for healthcare, advocating for models that harmonize performance and ethical responsibility with limited resources. Future efforts must prioritize clinician-AI collaboration and regulatory alignment to unlock scalable, human-centric solutions for NLP-centered clinical applications in English speaking territories. Although LLMs possess extensive medical knowledge, smaller, specialized clinical LLMs outperform them, even when fine-tuned on limited annotated data (Zhou et al., 2023b). However, generic SLM have potential to achieve comparable performance to LLMs in targeted use cases while reducing computational costs, emphasizing the sustainable model development (Allal et al., 2025). However, whether task-specific SLM in healthcare achieve comparable performance to LLMs or not, is still an open research question (Labrak et al., 2024; Fu et al., 2024; Zhang et al., 2025b). Efforts have been made to explore medical QA, language understanding and mental health analysis; a huge gap lies in model development for clinical concept extraction and report summarization. Current research on SLMs remains nascent, with limited exploration of longitudinal real-world deployments.
While this study focuses on NLP-centered healthcare informatics for English speaking territories, we acknowledge the growing importance of multilingual and non-English models, which is beyond our current scope. Our review excludes multimodal research and the model lacking peer-reviewed evidence. Moving forward, we aim to maintain an up-to-date repository on GitHub to reflect ongoing developments in this evolving field.

### v2

@@ 5.1. Current state-of-the-art
Discussion on comparative analysis of SLMs: SLMs have rapidly evolved for a diverse range of healthcare applications. Table 6 summarizes existing healthcare SLMs, detailing their base architectures, adaptation techniques, and target NLP tasks. Most SLMs are built on lightweight architectures like GPT-2, Phi-2, Mamba, Qwen, and LLaMA-2, often adapted with techniques such as fine-tuning and in-context learning. However, fewer models leverage quantization. Models like BioMistral, BioMistral-NLU, MentalQLM, and PMC-LLAMA exemplify state-of-the-art performance through multi-task adaptability, multilingual benchmarking, and specialized reasoning tasks (e.g., CoT QA). Yet, many efforts are still concentrated on QA and CLS, while areas like clinical SUM, temporal reasoning, and wellness dimension extraction remain underexplored. Additionally, we marked the availability (A) of SLMs in healthcare domain.
At present, healthcare models employ knowledge-aware methods, including ontology grounding using UMLS or SNOMED CT and RAG to enhance factual correctness. Integrating advanced interpretable reasoning mechanisms, such as metacognition, holds promise for enhancing the trustworthiness and reliability of CDSS. There is a pressing need to broaden the scope of tasks beyond traditional disease-centric applications like question answering and classification, toward holistic patient care, including the extraction and interpretation of emotional, social, behavioral, and wellness dimensions. Establishing standardized benchmarks for emerging clinical NLP tasks - wellness assessment, mental health triage, and personalized care planning will be vital to evaluating real-world utility and scientific progress. Moreover, Magnini et al. (2025) evaluated the use of locally deployed open-source generic SLMs for medical chatbots, focusing on hypertension self-management to ensure privacy and reliability. It demonstrates that lightweight SLMs can perform competitively in key tasks, offering a viable, privacy-preserving alternative to large proprietary models.
@@ 5.2. The role of SLM in healthcare
Beyond the performance of standalone models, we explore the role of SLMs in synergy with LLMs, emphasizing their complementary strengths, capabilities, and applications in healthcare research. The potential of collaborative and competitive roles of SLMs and LLMs is underappreciated in fostering efficient and accessible AI models for CDSS. Amid the broad spectrum of SLM applications observed in prior studies on general models (Chen and Varoquaux, 2024), this section delineates their specific relevance to NLP-driven decision-making tasks within the scope of healthcare informatics.
@@ 5.3. Assessment methods
Accuracy and Judgment: Chen et al. (2023) developed a novel, physician-curated adversarial question dataset rooted in real-world clinical scenarios, along with a comprehensive 17-metric evaluation framework designed to assess alignment and contextual relevance to clinical practice. Selene Mini is an open-source SLM designed for high-accuracy evaluation across diverse tasks [^14] https://huggingface.co/blog/AtlaAI/selene-1-mini. Trained with a principled blend of supervised and preference optimization on curated data, it outperforms 11 benchmarks of SLMs as a judge, demonstrating robust, domain-adaptable, and format-resistant performance for real-world deployment.
@@
Trustworthiness: Benchmarking LLMs in healthcare requires standardized, domain-specific metrics to ensure safety, accuracy, and ethical integrity while addressing data privacy, bias, and explainability (Budler et al., 2025; Aljohani et al., 2025). Truthfulness is critical in language models to prevent misinformation. Using the TruthfulQA, BioMistral-7B shows strong performance, outperforming GPT-3.5 Turbo by 4%, though no model excels across all categories, highlighting task-specific strengths and limitations (Labrak et al., 2024). Faithfulness in AI-generated medical information remains a critical concern, and recent efforts focus on understanding its causes, developing evaluation metrics, and proposing mitigation strategies to ensure factual accuracy across generative medical AI applications (Xie et al., 2023b).
Memory and energy efficiency: Energy efficiency is a high priority for SLMs, especially when used on battery-powered devices. MELODI is a novel framework and dataset designed to monitor and analyze energy consumption during LLM inference, revealing significant efficiency disparities and offering a foundation for sustainable, energy-conscious AI deployment (Husom et al., 2024). Response length significantly impact power consumption, and concise outputs can reduce energy usage. We measure the reduced environmental impact of healthcare SLM development through carbon footprints in Table 5.3. Greenhouse equivalences can be calculated with a well-establish calculator [^16] https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator.
@@ 6. Discussion and conclusion
Our goal is to provide healthcare professionals with practical insights on deploying SLMs in real-time applications, rather than explaining their internal workings. SLMs assist healthcare professionals by extracting key clinical concepts—such as SDoH, QoL, and therapeutic interventions—from unstructured text. They support CDSS through automated diagnosis assistance, personalized treatment recommendations, and summarization of lengthy patient records. Additionally, SLMs enhance real-time CDSS by identifying rare diseases, drug interactions, and aligning patient data with medical literature. SLMs support clinical research by structuring EHR data and monitoring adverse events, thereby reducing manual chart reviews and enhancing data usability. They automate clinical coding and billing by mapping free-text entries to standardized codes, minimizing administrative burden and claim errors. SLMs also enable clinical QAs by extracting relevant insights from research literature and patient records, accelerating evidence-based investigations. SLMs have the potential to enhance patient-centered care by analyzing sentiment and monitoring mental health to detect risks like worsening depression. They empower virtual health assistants and triage systems, facilitating personalized engagement and real-time conversational support. Additionally, SLMs improve healthcare accessibility through Plain Language Adaptation - simplifying complex medical information for individuals with limited health literacy.
@@ 6.1. Challenges and research directions
The contrast between the theoretical capabilities and real world limitations is a fascinating space to explore. SLMs in NLP-driven healthcare informatics are facing a significant benchmarking challenge. Unlike general-domain NLP, healthcare informatics relies heavily on proprietary, heterogeneous data, limiting reproducibility and transparency. Public datasets offer partial solutions but often lack representativeness. As models grow in complexity and autonomy, evaluation frameworks must evolve to incorporate safety, ethics, and trustworthiness. Standardized benchmarks are essential to ensure consistent assessment, with key criteria including “doctor-like” communication—demonstrating medical expertise, sound reasoning, and clinically relevant diagnostic insight. Progress requires collaborative efforts, sustainable funding, regulatory alignment, and a workforce skilled in both informatics and health sciences. Edge deployment is critical for expanding access to AI tools in underserved regions, enabling point-of-care decision support, and reducing reliance on cloud-based APIs that may raise data privacy concerns. Interoperability is the model’s ability to seamlessly integrate and operate within the existing Health Information Exchanges (HIEs) from different sites. Another significant advantage of SLMs is their potential when used in multi-agent AI systems—collections of specialized agents that collaborate to solve complex tasks. Instead of relying on a single monolithic model like an LLM, multi-agent systems distribute tasks among several domain-specific SLMs. Future research must focus on developing modular interfaces, shared embedding spaces, and universal chaining frameworks that enable plug-and-play interoperability across models of varying sizes and capabilities. We posit the SLM development as a multi-dimensional solution for advancing NLP-centered healthcare informatics. Several active research directions in the development of generic SLMs are being effectively adapted to advance healthcare-specific SLM development. We enlist significant developments in healthcare as follows.
@@
Human-in-the-loop systems integrate SLMs, LLMs, and expert oversight to ensure safe and effective deployment. SLMs assist with triage and uncertainty flagging, while LLMs manage complex synthesis. Embedding human feedback enhances trust and supports continual model refinement. Advancing flexible architectures, evaluation frameworks, and regulatory standards is essential, with future research focusing on efficient evaluators for clinical tasks, emphasizing factuality, safety, and uncertainty (Han et al., 2024; Xie et al., 2023b; Budler et al., 2025). Reducing manual chart review. As LLMs are preferentially aligned with human-preferred responses and surpass human capabilities, the feasibility and reliability of human evaluation diminish, introducing a weak-to-strong generalization, wherein smaller, less capable models supervise and guide the alignment of more powerful models (Burns et al., 2023). This paradigm has potential to reduce manual chart review efforts, enabling scalable, cost-effective, safe and factual preference alignment that exceed clinician expertise. Summarizing medical reports. SLMs perform high-quality abstractive summarization by distilling structured rationales from LLMs (Wang et al., 2025a). These rationales encapsulate important structural and semantic cues from the document, broken into high-level concepts, text-spans that support those themes, and relationships among key entities in the document. These structured rationales serve as interpretable reasoning traces that guide summary generation, similar to CoT prompting but in a structured, modular format. SLMs are rationale-aware, lightweight summarizers that are trained using intermediate reasoning signals setting a new paradigm for transparent and cost-effective summarization. Bias detection and de-biasing models. Because SLMs are often fine-tuned on relatively small clinical corpora (e.g., institutional EHRs or specialty-specific datasets), they are at high risk of inheriting data biases or overfitting to local linguistic patterns (Sapkota et al., 2025). This can manifest in poor generalization to diverse patient populations, inaccurate risk stratification, or propagation of health disparities in AI-driven decision-making. Future research must incorporate domain-aware debiasing techniques such as adversarial training, fairness-aware objective functions, and counterfactual data augmentation. Rigorous bias audits and alignment assessments should become standard in the development lifecycle of healthcare SLMs. Automated Machine Learning (AutoML). While existing model compression techniques have shown significant progress, they continue to rely heavily on manual design—particularly in defining suitable student architectures for knowledge distillation, which demands considerable human effort. The application of AutoML to LLM remains limited due to the significant computational burden involved in scaling down models with billions of parameters. SLMs address this challenge by evaluating architectural components like model depth and attention head width within a reduced parameter space (Liu et al., 2024e). Shen et al. (2024) reduced the search complexity by incorporating optimized initialization strategies, accelerating convergence. Therefore, the models must integrate AutoML approaches—meta-learning and neural architecture search, that offer a promising direction to mitigate this dependency. This will enhance compression efficiency without compromising performance.
@@ 6.2. Impact of SLMs
SLMs offer infrastructure-friendly deployment on edge devices using optimized hardware such as NVIDIA Jetson or Qualcomm AI accelerators. Software optimization techniques (distillation, quantization and pruning) significantly lower computational demands, enabling model adoption without costly hardware investments. However, challenges such as frequent model updates and deployment across heterogeneous devices may introduce new scalability and maintenance complexities. Given these considerations, investigating the impact of model development is necessary to support responsible application. Environmental and economic trade-offs: Training GPT-3 with 175 billion parameters consumed 1287 MWh of electricity, and resulted in carbon emissions of 502 metric tons of carbon, equivalent to driving 112 gasoline powered cars for a year (Patterson et al., 2021). The environmental impact of AI is an escalating concern (Rillig et al., 2023) where SLMs offer a partial solution by reducing the training and inference energy consumption by up to 80% (Liu and Yin, 2024). This reduction directly translates into a smaller carbon footprint, making SLMs a more sustainable choice as evident from Table 5.3. This efficiency is valuable in healthcare, aligning with the growing movement toward green AI, where the need of real-time solutions must be balanced with sustainability goals. The future directions for environmentally conscious green (small/large) language model development in healthcare promotes compression techniques discussed in §4. Ethical and Social Impact. The integration of SLMs into clinical workflows alleviates provider burden while promoting equitable access, reducing digital disparities, and enhancing patient-centered care. Their lightweight nature enables deployment in low-resource settings, including rural clinics, mobile health units, and edge devices, thereby extending AI-driven support to underserved populations. By lowering the technical and financial barriers associated with large models, SLMs democratize the use of advanced NLP tools across diverse healthcare institutions (Mittal et al., 2024). Additionally, SLMs support data de-identification by enabling efficient, on-device removal of PHI from clinical texts, facilitating privacy-preserving data sharing and compliance with healthcare regulations. SLMs can support patient friendly adoption through plain language adaptation, and empower patients with timely, comprehensible information. However, addressing fairness, bias, and equitable access remains critical (Sapkota et al., 2025). While the narrower focus of model training datasets allows for more effective bias filtering, it also risks excluding underrepresented groups or contexts. Therefore, careful dataset design is essential to ensure inclusive and responsible AI deployment. Synthesis and Perspectives. The goal of SLMs is to prioritize efficiency and accessibility. While, they mitigate environmental, economic, and ethical challenges associated with LLMs, the success of these methods largely depend on interdisciplinary collaborations, for example, hardware engineers optimizing for energy efficiency, clinicians validating medical tools, and policymakers ensuring equitable access. Future work in this direction requires standardized benchmarks for fairness and environmental impact, a step toward sustainable, human-centric AI.
@@ 6.3. Concluding remarks
This work underscores the transformative potential of SLMs in democratizing AI for healthcare, advocating for models that harmonize performance and ethical responsibility with limited resources. Future efforts must prioritize clinician-AI collaboration and regulatory alignment to unlock scalable, human-centric solutions for NLP-centered clinical applications in English speaking territories. Although LLMs possess extensive medical knowledge, smaller, specialized clinical LLMs outperform them, even when fine-tuned on limited annotated data (Zhou et al., 2023b). However, generic SLM have potential to achieve comparable performance to LLMs in targeted use cases while reducing computational costs, emphasizing the sustainable model development (Allal et al., 2025). Whether task-specific SLM in healthcare achieve comparable performance to LLMs or not, is still an open research question (Labrak et al., 2024; Fu et al., 2024; Zhang et al., 2025b). Efforts have been made to explore medical QA, language understanding and mental health analysis; a huge gap lies in model development for clinical concept extraction and report summarization. Current research on SLMs remains nascent, with limited exploration of longitudinal real-world deployments.
While this study focuses on NLP-centered healthcare informatics for English speaking territories, we acknowledge the growing importance of multilingual and non-English models, which is beyond our current scope. Our review excludes multimodal research and the SLMs lacking peer-reviewed evidence. Moving forward, we aim to maintain an up-to-date repository on GitHub to reflect ongoing developments in this evolving field.